# Reliable uncertainty with cheaper neural network ensembles: a case study in industrial parts classification

### v1

@@ Abstract
In operations research (OR), predictive models often encounter out-of-distribution (OOD) scenarios where the data distribution differs from the training data distribution. In recent years, neural networks (NNs) are gaining traction in OR for their exceptional performance in fields such as image classification. However, NNs tend to make confident yet incorrect predictions when confronted with OOD data. Uncertainty estimation offers a solution to overconfident models, communicating when the output should (not) be trusted. Hence, reliable uncertainty quantification in NNs is crucial in the OR domain. Deep ensembles, composed of multiple independent NNs, have emerged as a promising approach, offering not only strong predictive accuracy but also reliable uncertainty estimation. However, their deployment is challenging due to substantial computational demands. Recent fundamental research has proposed more efficient NN ensembles, namely the snapshot, batch, and multi-input multi-output ensemble. This study is the first to provide a comprehensive comparison of a single NN, a deep ensemble, and the three efficient NN ensembles. In addition, we propose a Diversity Quality metric to quantify the ensembles’ performance on the in-distribution and OOD sets in one single metric. The OR case study discusses industrial parts classification to identify and manage spare parts, important for timely maintenance of industrial plants. The results highlight the batch ensemble as a cost-effective and competitive alternative to the deep ensemble. It outperforms the deep ensemble in both uncertainty and accuracy while exhibiting a training time speedup of 7x, a test time speedup of 8x, and 9x memory savings.
@@ 1 Introduction
During deployment in operations research (OR), predictive models inevitably encounter scenarios where the data distribution differs from the training data distribution, referred to as out-of-distribution (OOD) data. For instance, in predictive maintenance, unseen operating conditions like higher temperatures or humidity might impact model performance.
In recent years, neural networks (NNs) are rapidly gaining traction in the OR field for their exceptional performance in fields such as image recognition (Madhav et al., 2023; Mena et al., 2023). However, when confronted with OOD data, NNs tend to make confident yet incorrect predictions (Guo et al., 2017). Specifically, an NN classifier may predict an incorrect label despite assigning a high predicted probability to that class. This behavior poses concerns as the model’s performance deteriorates over time with more OOD scenarios encountered, without warning the decision maker.
Uncertainty estimation provides a solution to overconfident models by capturing the uncertainty in both the data and the model. As such, it communicates when a model’s output should (not) be trusted (Thuy and Benoit, 2023). Therefore, reliable uncertainty quantification is crucial for the safe deployment of NNs in the OR domain.
Deep ensembles, comprising multiple independent NNs, have emerged as a promising approach, offering not only strong predictive accuracy but also reliable uncertainty estimation (Ovadia et al., 2019). Individual members are trained independently on the entire dataset and diversity arises from random weight initializations and sampling mini-batches of data during training. Diversity represents the level of disagreement among ensemble members. In OOD scenarios, deep ensembles raise warnings through increased uncertainty and diversity, facilitating timely human intervention to prevent misclassifications (Thuy and Benoit, 2023).
Despite their strong performance, deploying deep ensembles in practice is challenging due to the significant computational and memory demands. That is, the demands increase linearly with the ensemble size. Computation-wise, each member requires a separate forward pass during training and testing. Memory-wise, each member requires a copy of the NN weights. Allocating such substantial resources causes excessively long runtimes and is hard to justify considering financial constraints and environmental concerns.
Recent fundamental research has proposed more efficient NN ensembles, namely the snapshot ensemble (Huang et al., 2017), batch ensemble (Wen et al., 2020), and multi-input multi-output (MIMO) ensemble (Havasi et al., 2020). This study is the first to provide a comprehensive comparison of a single NN, a deep ensemble, and the three efficient NN ensembles. In addition, we propose a Diversity Quality ($DQ_{1}$) score to quantify the ensembles’ performance on the in-distribution (ID) and OOD sets in one single metric, as opposed to two separate metrics in current literature. The OR case study discusses industrial parts classification to identify and manage spare parts. This task is crucial for timely maintenance of industrial plants and to avoid downtime. During deployment, classifiers frequently encounter unknown parts not present in the training set, requiring a prompt for human intervention.
@@
Extensive analysis of the case study results illustrates that batch ensemble offers a robust alternative to deep ensemble, delivering superior performance at a fraction of the computational cost. In contrast, snapshot and MIMO ensemble lag behind due to either low predictive performance or poor diversity quality.
@@
The remainder of the paper is organized as follows. Section 2 gives an overview of related work and section 3 presents the efficient ensembling techniques. In Section 4, the case study in industrial parts classification is presented. Section 5 gives the results and discussion and presents the Diversity Quality score. Finally, section 6 provides a conclusion.
@@ 2 Related Work
Related work in OR uses NN ensembles to either improve the predictive performance or to quantify uncertainty. This section describes the literature in both streams.
2.1 Ensembles for improved predictive performance
Ensembles have been primarily used to improve the predictive performance of a machine learning system. The individual ensemble members are combined by averaging or majority voting the predictions of all members. Averaging over $M$ members goes as follows:
@@ 2.1 Ensembles for improved predictive performance
NNs have been widely used as an ensemble member in heterogeneous ensembles, together with other machine learning models such as Lasso regression, XGBoost, and Support Vector Machines. Literature ranges over multiple fields, from finance (Du Jardin, 2021; Cui et al., 2022; Zhang et al., 2023; Jiang et al., 2022; Zhang et al., 2022; Li and Chen, 2021; Krauss et al., 2017) over healthcare (Abedin et al., 2021; Baradaran Rezaei et al., 2023) and manufacturing (Yang et al., 2021; Wu et al., 2022), to politics (Easaw et al., 2023).
Homogeneous ensembles of NNs have also been applied in OR. Relevant literature is in customer service (Zicari et al., 2022), healthcare (Poloni et al., 2022), project tendering (Bilal and Oyedele, 2020), manufacturing (Gupta et al., 2023), and tourism (Pitakaso et al., 2023). The studies find that by combining the outputs of several NNs, an ensemble can outperform any of its members. However, this body of literature is smaller than heterogeneous ensembles, possibly due to the high computational overhead of training multiple NNs.
2.2 Ensembles for improved uncertainty quantification
In addition to having strong predictive performance, deep Ensembles have also been shown to exhibit reliable uncertainty estimates (Lakshminarayanan et al., 2017). Uncertainty arises from two sources: data and model uncertainty (Der Kiureghian and Ditlevsen, 2009). Data uncertainty refers to the notion of randomness and is related to the data-measurement process. This uncertainty is irreducible even if more data is collected. Model uncertainty accounts for uncertainty in the model parameters, i.e., uncertainty about which model generated the collected data. In contrast to data uncertainty, collecting more data can reduce model uncertainty. Both types of uncertainty can then be summed to compute the total uncertainty in a prediction.
@@ 2.2 Ensembles for improved uncertainty quantification
First, total uncertainty and data uncertainty are approximated using classical information-theoretic measures; then model uncertainty is obtained as the difference (Depeweg et al., 2018):
@@
$u_{total}(\mathbf{x}^{*})$ $\approx H\left[\frac{1}{M}\sum_{m=1}^{M}p(\mathbf{y}^{*}\mid\mathbf{x}^{*},\ \bm{\theta}_{m})\right]$ (2)
@@
$u_{data}(\mathbf{x}^{*})$ $\approx\frac{1}{M}\sum_{m=1}^{M}H\left[p(\mathbf{y}^{*}\mid\mathbf{x}^{*},\ \bm{\theta}_{m})\right]$ (3)
$u_{model}(\mathbf{x}^{*})$ $=u_{total}(\mathbf{x}^{*})-u_{data}(\mathbf{x}^{*}).$ (4)
@@
Total uncertainty is computed by averaging the predictive distributions over the different members and calculating the entropy $H$. Data uncertainty is computed by calculating the entropy in each member and averaging the entropies. Intuitively, data uncertainty measures uncertainty in the softmax classification on individual members; model uncertainty measures how much the members deviate. Naturally, the single NN has zero model uncertainty by definition because it only consists of one member. For a classification problem with $K$ classes, the maximum uncertainty in a prediction is $\log(K)$.
Uncertainty quantification is underinvestigated in OR with a limited body of literature. Deep ensembles have been used by Thuy and Benoit (2023); Han and Li (2022); Kim et al. (2023); Prasad et al. (2024); Wen et al. (2022) in the context of OOD detection. Predictions with high uncertainty indicate that the observations lie outside the training data distribution. As such, these predictions are discarded and the observations are passed on to a human expert for a label. Note that Wen et al. (2022) use a snapshot ensemble, although its performance is not compared to other ensemble techniques or a single NN. Furthermore, Zou and Chen (2021) use deep ensembles in active learning. Observations in the unlabeled set causing high model uncertainty are interesting to label. That is, these observation lie outside the current training data distribution and can reduce the model uncertainty once added to the labeled training set.
Kraus et al. (2020) report that the lack of uncertainty quantification in single NNs is a key limiting factor for adoption in the field of OR. The reliable uncertainty estimates from deep ensembles could increase the relevance of NNs, highlighting the need for more efficient NN ensembles.
@@ 3.1 Snapshot ensemble
The maximum learning rate controls how far the optimizer jumps to another mode after a restart. Note that the learning rate schedule is determined by the procedure and cannot be combined with another custom learning rate schedule. The selected number of ensemble members directly affects the training time of each member; too many members damages their individual accuracy while too few members negatively affects the uncertainty estimates.
The snapshot ensemble has the same training time than a single NN. However, similar to deep ensembles, the memory cost and the computational cost at test time increase linearly with the amount of ensemble members.
@@ 3.2 Batch ensemble
The fast weights can be initialized with either a random signed vector or a random Gaussian vector. Wen et al. (2020) note that the random signed vector often results in higher diversity among the members. In theory, more ensemble members give better results but a higher computational cost; in this sense it is similar to deep ensemble.
@@ 3.3 Multi-Input multi-Output (MIMO) ensemble
The MIMO configuration with $M$ members requires two changes to the NN architecture (Figure 3). First, the input layer is replaced, taking $M$ datapoints instead of a single datapoint. Second, the output layer is replaced, consisting of $M$ classification heads based on the last hidden layer instead of a single head. During training, the $M$ inputs are sampled independently from the training set and each of the $M$ heads are trained to predict its matching input. At test time, the same input is repeated $M$ times, causing the heads to make $M$ independent predictions on the same input, effectively forming an ensemble.
@@
The number of ensemble members directly affects the number of relationships that need to be learned, which determines the capacity of the subnetworks. Too many members might make the NN capacity prohibitively small with poor accuracy as a result; too few members will cause poor diversity. Havasi et al. (2020) report that 3–4 ensemble members is typically optimal. Additionally, the independence between inputs can be relaxed so that subnetworks can share features. Relaxing independence might particularly benefit networks with limited excess capacity. Increasing the input repetition probability improves the predictive performance but negatively impacts the diversity.
@@
4 Case Study: Industrial Parts Classification
4.1 Problem setting
Industrial parts classification is crucial for inventory control of spare parts and for maintaining the smooth operation of industrial plants. Spare parts, essential for timely maintenance of industrial plants, constitute a significant portion of inventory stocks (Hu et al., 2017). According to Hu et al. (2015), operational and maintenance support expenses often represent over 60% of the total costs in industrial settings, with spare parts alone accounting for 25–30%. In large enterprises, the inventory can reach tens of thousands of items (Ernst and Cohen, 1990), making it impractical for human operators to manually identify each part and its appropriate inventory strategy. Industrial parts classification offers a solution by enabling the identification and management of spare parts efficiently.
During deployment, a parts classifier is likely to encounter unknown parts not present in the training data. This might occur, for instance, if new parts are introduced after the classifier’s training phase or if it encounters parts handled by departments not included in the initial training data. It is imperative that the classifier can identify these novel situations and signal a warning, prompting human intervention, rather than making incorrect predictions with high confidence.
We apply the efficient ensembling techniques to a dataset for industrial parts classification, investigating their predictive performance and uncertainty estimates, along with their computational expense. Furthermore, we compare the results to those obtained from a single NN and a deep ensemble. The models are trained and subsequently evaluated on (i) an ID test set comprising the same objects and (ii) an OOD set containing entirely different objects.

### v2

# Fast and reliable uncertainty quantification with neural network ensembles for industrial image classification

@@ Abstract
Image classification with neural networks (NNs) is widely used in industrial processes, situations where the model likely encounters unknown objects during deployment, i.e., out-of-distribution (OOD) data. Worryingly, NNs tend to make confident yet incorrect predictions when confronted with OOD data. To increase the models’ reliability, they should quantify the uncertainty in their own predictions, communicating when the output should (not) be trusted. Deep ensembles, composed of multiple independent NNs, have been shown to perform strongly but are computationally expensive. Recent research has proposed more efficient NN ensembles, namely the snapshot, batch, and multi-input multi-output ensemble. This study investigates the predictive and uncertainty performance of efficient NN ensembles in the context of image classification for industrial processes. It is the first to provide a comprehensive comparison and it proposes a novel Diversity Quality metric to quantify the ensembles’ performance on the in-distribution and OOD sets in one single metric. The results highlight the batch ensemble as a cost-effective and competitive alternative to the deep ensemble. It matches the deep ensemble in both uncertainty and accuracy while exhibiting considerable savings in training time, test time, and memory storage.
@@ 1 Introduction
In recent years, neural networks (NNs) have rapidly gained traction in operations research (OR) (Mena et al., 2023) and have demonstrated strong performance in handling high-dimensional inputs in image classification for industrial settings (Madhav et al., 2023). During deployment, image classifiers inevitably encounter scenarios where the data distribution differs from the training data distribution. It may specifically find unseen operating conditions, referred to as shifted data or covariate shift (Quiñonero-Candela et al., 2022), and unseen objects, referred to as entirely out-of-distribution (OOD) data. For instance, in industrial parts classification, images of known objects in unseen positions or lighting conditions are shifted data; newly developed objects are OOD data. Furthermore, for example in defect detection, unseen defect types represent OOD data. However, when confronted with such shifted or OOD data, NNs tend to make confident yet incorrect predictions (Guo et al., 2017). Specifically, an NN classifier may predict an incorrect label despite assigning a high predicted probability to that class. This behavior is problematic because the model’s performance deteriorates over time without warning the practitioner.
To ensure their reliability, NNs should quantify the uncertainty in their own predictions, referred to as model or epistemic uncertainty. This notion of epistemic uncertainty is not captured by standard NN classifiers and communicates when an NN’s output should (not) be trusted, crucial for the safe deployment of NNs in the OR domain (Thuy and Benoit, 2023). In this respect, deep ensembles, comprising multiple independent NNs, have emerged as a popular approach to quantify epistemic uncertainty. Under shifted and OOD data, deep ensembles raise warnings through increased uncertainty, facilitating timely human intervention to prevent misclassifications (Ovadia et al., 2019; Thuy and Benoit, 2023).
Despite their strong performance, deploying these ensembles in practice is challenging due to the significant computational and memory demands. First, low inference times are required for a classification system to be relevant in practice. For example in a sorting station, the predictions of a parts classifier should be available quickly enough to keep up with the speed of the conveyor belt. Second, long training times are impractical and allocating substantial resources is hard to justify considering financial constraints and environmental concerns. Third, limited memory requirements are important when deploying the NN on-device with restricted compute power. For example, a mobile device with a camera installed over a sorting station should have sufficient compute to perform inference. Recent fundamental research has proposed more efficient NN ensembles, namely the snapshot ensemble (Huang et al., 2017), batch ensemble (Wen et al., 2020), and multi-input multi-output (MIMO) ensemble (Havasi et al., 2020).
This study investigates the predictive and uncertainty performance of efficient NN ensembles in the context of image classification for industrial processes. The experiments are performed on the SIP-17 dataset (Zhu et al., 2023), a dataset of synthetic images for industrial parts classification. The contributions are twofold: (i) it is the first to provide a comprehensive comparison of a single NN, a deep ensemble, and the three efficient NN ensembles; (ii) we propose a Diversity Quality ($DQ_{1}$) score to quantify the ensembles’ performance on the in-distribution (ID) and OOD sets in one single metric, as opposed to two separate metrics in current literature.
@@
Experimental analysis illustrates that batch ensemble offers a robust alternative to deep ensemble, delivering similar performance at a fraction of the computational cost. In contrast, snapshot and MIMO ensemble lag behind due to either low predictive performance or poor Diversity Quality.
@@
The remainder of the paper is organized as follows. Section 2 gives an overview of related work, Section 3 discusses uncertainty estimation with ensembles and Section 4 presents the efficient ensembling techniques. In Section 5, the experiments in industrial parts classification are presented and the Diversity Quality score is outlined. Section 6 gives the results and discussion. Finally, Section 7 provides a conclusion.
@@ 2 Related Work
NN ensembles are commonly used in OR studies to either improve predictive performance or obtain reliable uncertainty estimates.
For predictive performance, NNs are most often used as an ensemble member in heterogeneous ensembles, together with other machine learning models such as Lasso regression, XGBoost, and Support Vector Machines. Literature ranges over multiple fields, from manufacturing (Yang et al., 2021; Wu et al., 2022) over healthcare (Abedin et al., 2021; Baradaran Rezaei et al., 2023) and finance (Du Jardin, 2021; Cui et al., 2022; Zhang et al., 2023; Jiang et al., 2022; Zhang et al., 2022; Li and Chen, 2021; Krauss et al., 2017), to politics (Easaw et al., 2023). Furthermore, although less common than heterogeneous ensembles of NNs and machine learning models, some studies use homogeneous ensembles of NNs, i.e. deep ensembles. Relevant literature on deep ensembles is in manufacturing (Gupta et al., 2023), project tendering (Bilal and Oyedele, 2020), customer service (Zicari et al., 2022), healthcare (Poloni et al., 2022), and tourism (Pitakaso et al., 2023). The studies find that by combining the outputs of several NNs, an ensemble can outperform any of its members.
Despite the extensive use of NN ensembles for predictive performance, uncertainty quantification for NNs is underinvestigated in OR. Deep ensembles have been used by Thuy and Benoit (2023); Han and Li (2022); Kim et al. (2023); Prasad et al. (2024); Wen et al. (2022) in the context of OOD detection, reducing misclassifications using classification with rejection (Mena et al., 2021). In this workflow, uncertain predictions are discarded and the observations are passed on to a human expert for a label. Note that Wen et al. (2022) employ a type of snapshot ensemble, although its behavior is not compared to other ensemble techniques or a single NN. Furthermore, Zou and Chen (2021) use deep ensembles in active learning, guiding labeling efforts based on epistemic uncertainty estimates. Heterogeneous ensembles have not been used for uncertainty quantification.
Overall, deep ensembles are appealing to practitioners. Compared to heterogeneous ensembles of NNs and other machine learning models, deep ensembles are more computationally expensive but easier to implement, as only one underlying model needs to be developed. Furthermore, compared to heterogeneous ensembles of different NN architectures, deep ensembles are computationally cheaper and easier to implement.
Furthermore, classification with rejection based on uncertainty quantification from deep ensembles is preferred over separate rejector-predictor architectures. That is, such an architecture first uses an anomaly detection algorithm to decide whether to reject the observation; if the observation is not rejected, the predictor is used for inference (Hendrickx et al., 2024). A separated rejector is usually simpler to operationalize but often results in sub-optimal rejection performance as there is no information sharing between the rejector and predictor (Homenda et al., 2014). Integrated rejection based on NN uncertainty estimates has the additional benefit that the predictor is better, as a deep ensemble typically has stronger predictive performance than a single NN.
Kraus et al. (2020) report that the lack of epistemic uncertainty quantification in single NNs is a key limiting factor for adoption in the field of OR. The reliable uncertainty estimates from deep ensembles could increase the relevance of NNs, highlighting the need for more efficient NN ensembles that bring this uncertainty information at a lower computational cost.
3 Uncertainty Quantification with Ensembles
3.1 Aleatoric and epistemic uncertainty
Uncertainty arises from two sources: aleatoric and epistemic uncertainty (Der Kiureghian and Ditlevsen, 2009). Aleatoric uncertainty refers to the notion of randomness and is related to the data-measurement process. This uncertainty is irreducible even if more data is collected. Epistemic uncertainty accounts for uncertainty in the model parameters and is naturally high for observations outside the training data distribution. In contrast to aleatoric uncertainty, collecting more data can reduce epistemic uncertainty. The sum of these two types represents the total uncertainty in a prediction.
NN classifiers capture aleatoric uncertainty in their softmax output distribution, forming a categorical distribution over the class labels. Capturing epistemic uncertainty requires an ensemble of NN classifiers in order to evaluate the disagreement among the ensemble members. A naive ensemble of NNs, i.e., a deep ensemble, is an effective but expensive way to obtain reliable epistemic uncertainty estimates (Lakshminarayanan et al., 2017).
3.2 Deep ensembles
A deep ensemble (Lakshminarayanan et al., 2017) is an ensemble of $M$ individual NNs, with each member trained independently on the entire dataset using all input features. The diversity among the members arises from random weight initializations and from randomness in sampling mini-batches of data during training, resulting in different solutions throughout the complex loss landscape. At test time with test sample $\mathbf{x}^{*}$, each of the $M$ members performs one forward pass on the input. The resulting $M$ softmax distributions are averaged to obtain the ensemble prediction $p(\mathbf{y}^{*}\mid\mathbf{x}^{*})$:
@@ 3.2 Deep ensembles
Deep ensembles are computationally expensive because their demands increase linearly with the ensemble size $M$. Computation-wise, each member requires a separate forward pass during training and testing. Memory-wise, each member requires a copy of their NN weights. Lakshminarayanan et al. (2017) experiment with various ensemble sizes, ranging from 5 to 15 members. They find that a small ensemble with 5 members already improves significantly over a single NN, and that using more than 10 members results in limited improvement.
3.3 Uncertainty decomposition
Each NN ensemble, naive or efficient, has $M$ predictions for each test sample $\mathbf{x}^{*}$. Based on these predictions we can calculate the total uncertainty, which in turn can be further decomposed in aleatoric and epistemic uncertainty.
@@ 3.3 Uncertainty decomposition
First, total uncertainty $\textrm{TU}(\mathbf{x}^{*})$ and aleatoric uncertainty $\textrm{AU}(\mathbf{x}^{*})$ are approximated using classical information-theoretic measures; then epistemic uncertainty $\textrm{EU}(\mathbf{x}^{*})$ is obtained as the difference (Depeweg et al., 2018):
@@
$\textrm{TU}(\mathbf{x}^{*})$ $\approx\mathbb{H}\left[\frac{1}{M}\sum_{m=1}^{M}p(\mathbf{y}^{*}\mid\mathbf{x}^{*},\ \bm{\theta}_{m})\right]$ (2)
@@
$\textrm{AU}(\mathbf{x}^{*})$ $\approx\frac{1}{M}\sum_{m=1}^{M}\mathbb{H}\left[p(\mathbf{y}^{*}\mid\mathbf{x}^{*},\ \bm{\theta}_{m})\right]$ (3)
$\textrm{EU}(\mathbf{x}^{*})$ $=\textrm{TU}(\mathbf{x}^{*})-\textrm{AU}(\mathbf{x}^{*}).$ (4)
@@
Total uncertainty is computed by averaging the predictive distributions (i.e., softmax distributions) over the different members and calculating the entropy $\mathbb{H}$. Aleatoric uncertainty is computed by calculating the entropy in each member and averaging the entropies. Intuitively, aleatoric uncertainty measures uncertainty in the softmax output of individual members; epistemic uncertainty measures how much the members deviate. Naturally, the single NN has zero epistemic uncertainty because it only consists of one member. For a classification problem with $K$ classes, the maximum uncertainty in a prediction is $\log(K)$.
Table 1 contains three examples in the context of binary classification with $M=4$ ensemble members. The middle and bottom rows both have a total uncertainty of $1.0$ although the predictions are wildly different. Therefore, the total uncertainty alone is insufficient to characterize the NN’s predictions; decomposition into aleatoric and epistemic uncertainty is necessary.
Table 1: Examples of uncertainty decomposition (adapted from Thuy and Benoit (2023))
Predictions $p(\mathbf{y}^{*}\mid\mathbf{x}^{*},\ \bm{\theta}_{m})$ $p(\mathbf{y}^{*}\mid\mathbf{x}^{*})$ $\textrm{TU}(\mathbf{x}^{*})$ $\textrm{AU}(\mathbf{x}^{*})$ $\textrm{EU}(\mathbf{x}^{*})$
{(1.0, 0.0), (1.0, 0.0), (1.0, 0.0), (1.0, 0.0)} (1.0, 0.0) 0. 0. 0.
{(0.5, 0.5), (0.5, 0.5), (0.5, 0.5), (0.5, 0.5)} (0.5, 0.5) 1. 1. 0.
{(1.0, 0.0), (0.0, 1.0), (1.0, 0.0), (0.0, 1.0)} (0.5, 0.5) 1. 0. 1.
@@ 4.1 Snapshot ensemble
The maximum learning rate controls how far the optimizer jumps to another mode after a restart. Note that the learning rate schedule is determined by the procedure and cannot be combined with another custom learning rate schedule. The selected number of ensemble members directly affects the training time of each member; too many members damages their individual accuracy while too few members negatively affects the uncertainty estimates. Huang et al. (2017) find that ensemble sizes of 4–8 work well.
The snapshot ensemble has the same training time as a single NN. However, similar to deep ensembles, the memory cost and the computational cost at test time increase linearly with the number of ensemble members.
@@ 4.2 Batch ensemble
The fast weights can be initialized with either a random sign vector or a random Gaussian vector. Wen et al. (2020) note that the random signed vector often results in higher diversity among the members. The fast weight can also be updated differently than the slow weights using a learning rate multiplier smaller than 1. In theory, more ensemble members give better results but a higher computational cost; in this sense, it is similar to the deep ensemble. Wen et al. (2020) experiment with ensemble sizes 4 and 8 and find that size 8 generally performs best.
@@ 4.3 Multi-input multi-output (MIMO) ensemble
The MIMO configuration with $M$ members requires two changes to the NN architecture (Figure 3). First, the input layer is replaced, taking $M$ data points instead of a single data point. Second, the output layer is replaced, consisting of $M$ classification heads based on the last hidden layer instead of a single head. During training, the $M$ inputs are sampled independently from the training set and each of the $M$ heads is trained to predict its matching input. At test time, the same input is repeated $M$ times, causing the heads to make $M$ independent predictions on the same input, effectively forming an ensemble.
@@
The number of ensemble members directly affects the number of relationships that need to be learned, which determines the capacity of the subnetworks. Too many members might make the NN capacity prohibitively small with poor accuracy as a result; too few members will cause poor diversity. Havasi et al. (2020) report that 3–4 ensemble members is typically optimal. Additionally, the independence between inputs can be relaxed so that subnetworks can share features. Relaxing independence might particularly benefit networks with limited excess capacity. Increasing the input repetition probability improves the predictive performance but negatively impacts the diversity. Furthermore, examples can be repeated in the minibatch in order to achieve more stable gradient updates, which also has an implicit regularization effect.
@@
5 Experiments: Industrial Parts Classification